# Geometric-Aware Low-Light Image and Video Enhancement via Depth Guidance

### v1

@@ 4.3 Ablation Study
Methods PSNR SSIM PSNR SSIM
Ours with Decoder 28.97 0.875 27.34 0.799
Ours w/o Corre. 30.74 0.894 24.82 0.794
Ours with Add 30.57 0.892 27.14 0.786
Ours with $\lambda\times 5$ 29.53 0.878 27.99 0.802
Ours with $\lambda\div 5$ 29.89 0.883 27.42 0.795
Ours 30.95 0.897 27.89 0.804
Table 3: Results of the ablation study in the SDSD-indoor and SDSD-outdoor datasets with DP3DF. Our method performs better than all ablation settings.
SDSD-indoor SDSD-outdoor
Methods PSNR SSIM PSNR SSIM
Ours with Decoder 23.50 0.877 29.14 0.836
Ours w/o Corre. 22.86 0.876 27.33 0.823
Ours with Add 25.24 0.902 28.06 0.834
Ours with $\lambda\times 5$ 23.92 0.900 30.00 0.853
Ours with $\lambda\div 5$ 23.75 0.896 30.03 0.851
Ours 26.86 0.924 29.34 0.849
@@
Table 4: Results of the ablation study in the SDSD-indoor, SDSD-outdoor, and SMID datasets with SNR. Our method performs better than all ablation settings.
4.4 User Study
@@
For more details, we initially select $T$ videos/images randomly from each testing dataset for a given LLIE/LLVE approach, where $T=20$ for LLIE and $T=10$ for LLVE. Our user study employs an AB-test strategy, where each participant is presented with an image/video from the corresponding baseline and an image/video from our framework. Importantly, participants are unaware of which one is the baseline, as the order of our results and baseline results is randomized during each evaluation. During each evaluation, participants compare 5 pairs for a specific method on a given dataset. They have the option to choose which one is better or indicate that the two are the same. Participants are instructed to evaluate the results based on criteria such as natural color, minimal noise, realistic illumination, and stable temporal changes (for LLVE). Each participant is provided with an example. Each participant completes 15 tasks (10 methods × 5 pairs). It takes approximately 30 minutes for a participant to finish the user study.
@@
The results of the user study are presented in Fig. 6. Clearly, the participants consistently preferred the results of our method. This indicates that our approach significantly improves the human subjective perception of existing LLIE and LLVE frameworks.
Figure 6: The results of the user study, which summarize that the results enhanced with our strategy, are preferred by participants compared with the baselines’ results.
5 Limitation
@@
In this paper, we introduce a novel emphasis on employing geometrical priors, particularly in the form of depth, for low-light enhancement, a aspect overlooked by previous methods. Our extensive experiments across various datasets and networks demonstrate the effectiveness of our proposed HDGFFM. Despite achieving significant advancements, there are still areas for improvement. Firstly, further efforts are required to minimize model parameters and additional computational requirements of HDGFFM, even though it has not introduced heavy computations in this paper. Secondly, refining our frameworks to establish a greater distinction between ours and baselines is a future direction. This could be achieved by incorporating other 3D priors.
6 Conclusion
@@
This paper proposes the utilization of geometrical priors (depth in this paper) for low-light enhancement, encompassing both image and video enhancement. To distill depth estimation capabilities from large-scale open-world depth estimation models, we introduce a lightweight depth estimator. Furthermore, the extracted depth-aware features are efficiently fused with the encoder features of the target network using our proposed HDGFFM, which incorporates correlation-based cross attention for fusion. Through extensive experiments conducted on representative datasets and methods, we demonstrate the effectiveness of our proposed strategy.

### v2

@@ IV-B Comparison
Methods PSNR SSIM LPIPS RMSE PSNR SSIM LPIPS RMSE
KinD(Supervised)[43] 19.13 0.876 0.346 53.462 18.10 0.761 0.346 28.540
URetinex-Net++(Supervised)[54] 20.29 0.922 0.204 46.030 17.70 0.520 0.265 47.800
LLFlow-L-SKF(Supervised)[12] 21.58 0.980 0.275 40.260 24.89 0.872 0.423 22.085
Retinexmamba(Supervised)[45] 22.84 0.985 0.015 7.490 18.56 0.939 0.072 15.860
Wakeup-Darkness(unsupervised)[27] 18.51 0.714 0.413 46.032 18.43 0.710 0.365 36.256
Zero-DCE(Unsupervised)[46] 17.81 0.787 0.361 58.046 17.70 0.520 0.265 47.796
RUAS(Unsupervised)[47] 13.97 0.717 0.467 82.992 11.19 0.469 0.307 92.194
SCI(Unsupervised)[48] 17.54 0.872 0.389 60.102 21.68 0.783 0.283 32.570
NeRCO(Unsupervised)[55] 21.73 0.849 0.185 41.130 22.25 0.843 0.334 38.261
Clip-LIE(Unsupervised)[49] 16.72 0.852 0.393 64.417 20.59 0.862 0.318 34.492
Ours+SNR 26.86 0.924 0.063 13.028 29.38 0.858 0.135 9.429
TABLE X: The quantitative comparison illustrates the range of plausible outcomes for methods with and without our strategy. We report the average variance value here. Incorporating our strategy reduces the variance among outputs from different low-light enhancement methods under the same supervision, thereby alleviating the ill-posed nature of the problem.
Methods SDSD-indoor SDSD-outdoor SMID LOLv2-real LOLv2-synthetic SID
[…]
TABLE XI: The average $L_{2}$ distance between the features of low-light and normal-light images reflects their alignment potential, demonstrating the effectiveness of our learned features.
Methods SDSD-indoor SDSD-outdoor SMID LOLv2-real LOLv2-synthetic SID
[…]
Figure 4: Visual comparison on different datasets. From top to down, they are Restormer in SMID and SDSD-outdoor, and SNR in LOLv2-real. Our method yields better results.
Figure 5: Visual comparisons of Restormer in the SID dataset.
Figure 6: Visual comparisons of SNR in the LOLV2-synthetic dataset.
Figure 7: Visual comparisons of SNR in the SID dataset.
Evaluation for the range of plausible outcomes. As stated in the introduction, we think that incorporating priors imposes additional constraints on the enhancement process, thereby reducing the range of plausible outcomes. This statement highlights that low-light image enhancement is a highly ill-posed problem, where even under identical supervision (i.e., the same low-light and normal-light data pairs), the outputs from different methods can vary significantly while remaining plausible. To verify this claim, we conduct a corresponding evaluation. We compute the distances between outputs of different LLIE methods and the corresponding ground truths. These LLIE methods include Retinexformer, Restormer, and SNR. For each image in the evaluation datasets, this yields a triplet representing the variation in “output deviation from ground truth” achievable by LLIE models. The variance of these triplets reflects the range of plausible outcomes for the dataset, with lower variance indicating a narrower range. As shown in the Table XI, we perform this evaluation across different datasets and compute the average variance of the triplets for all images within each dataset. The results demonstrate that the range of plausible outcomes produced by our method with priors is significantly narrower than that of the baselines, supporting our claim.
Analyses: our features are indeed beneficial? We have designed an experiment to evaluate whether the deep features in our method are effectively benefited. Since our fusion of image features and depth-aware features occurs in the encoder, we focus on analyzing the features in the deepest layer. Specifically, we feed low-light data into the network to obtain “feature A” and normal-light data to obtain “feature B” (to this end, we finetune the network using normal-light input data with a reconstruction loss). The distance between these features reflects the distribution deviation between low-light and normal-light data in the network’s deep feature space. Since normal-light features are more easily aligned with the ground truth, a smaller distance between “feature A” and “feature B” indicates more beneficial features for improved alignment. As shown in the Table XI, our method yields smaller distances, demonstrating that the enhanced features are indeed beneficial.
Qualitative evaluation. We also provide the visual comparisons for the baseline and ours on different datasets, as shown in Figs. 4, 5, 6, 7, and 8. As we can see, our method’s results are closer to the ground truth with natural illumination and color, fewer noise and artifacts. Our approach offers several well-founded advantages: the depth information provides additional 3D cues, enabling us to capture the actual lighting conditions that conform to physical laws. Consequently, the depth features can contribute to color accuracy, sharper edge boundaries, and better foreground/background separation.
Figure 8: Visual comparisons of Retinexformer in the SID (top) and SDSD-outdoor dataset (bottom).
IV-C Ablation Study
In this section, we perform various ablation studies on the SDSD datasets, encompassing both indoor and outdoor subsets, to assess the significance of different components in our framework. Specifically, we evaluate the key components in HDGFFM through several kinds of ablated cases: (i) The location where the geometric factor is added. (ii) The method of fusion between image features and depth-aware features. (iii) Different weights for depth loss. (iv) The influence of the choice on pre-trained depth estimator. (v) Others.
  • $\bullet$
    “Ours with Fusion in Decoder”. As indicated in Sec. III-A and Eq 2, the fusion in our framework occurs in the encoder of the target network $\mathcal{G}$. This choice is made because both the image content features and depth-aware features are extracted from the low-light input image, fostering homogeneity and enhancing fusion effects. To investigate the impact of alternative fusion points, we aim to test whether fusing image content features and depth-aware features in the decoder of $\mathcal{G}$ yields comparable results.
  • “Ours w/o $w_{l}$. In this ablation setting, we omit the correlation computation step (i.e., $w_{l}$ in Eq 8), and the fusion procedure (Eq 9) can be written as $\bar{f}_{g,l}=f^{\prime\prime}_{g,l}+f^{\prime}_{g,l}+f_{g,l}$.
  • $\bullet$
    “Ours with Alternative Attention”. In this ablation study, we explore the impact of different attention map formulation strategies for acquiring information from depth-aware features. In other words, we compute the attention map using only the depth-aware features, and employ the original image features to predict the value part: replacing $W_{q}(f_{g,l})$ in Eq 7 as $\hat{W}_{q}(\hat{f}_{d,l})$ and $\hat{W}_{v}(\hat{f}_{d,l})$ as $\hat{W}_{v}(f_{g,l})$.
  • $\bullet$
    “Ours with $\lambda\times 5$” and “Ours with $\lambda\div 5$. As specified in Eq 12, there is a hyper-parameter that governs the contribution of image reconstruction loss and depth prediction loss, denoted as $\lambda$. We aim to analyze the effects of varying the value of $\lambda$ on the final results.
  • $\bullet$
    “Ours with Depth Anything” and “Ours with Depth Anything2”. We integrate a new SOTA depth estimation approach, Depth Anything [56] and Depth Anything V2 [57], into our framework to replace DPT, demonstrating that our method’s performance is not dependent on the choice of the pre-trained $\mathcal{D}$.
  • $\bullet$
    “Depth Feature I” and “Depth Feature II”. In this ablation setting, we study whether our depth-aware image feature is better than depth features extracted from depth maps. Thus, we set two ablation experiments. “Depth Feature I” means extracting depth features directly from low-light images’ depth maps; “Depth Feature II” means obtaining depth features from initially enhanced low-light images’ depth maps.
  • $\bullet$
    “NYU/KITTI finetune”. We aim to explore the impact of using datasets with real-captured depth, such as NYU or KITTI, and synthesizing corresponding low-light and normal-light pairs. The experimental results are obtained by finetuning a model with these additional NYU/KITTI samples.
  • $\bullet$
    “depth estimator params+”. We aim to analyze whether the performance of our framework can be enhanced with the enhanced model capability of the depth estimator $\mathcal{F}$. We conduct an experiment by increasing the model capability of the depth estimator, using a transformer.
@@ IV-C Ablation Study
TABLE XII: Results of the ablation study in the SDSD-indoor, SDSD-outdoor and SMID datasets with DP3DF. Our method performs better than all ablation settings.
SDSD-indoor SDSD-outdoor SMID
Methods PSNR SSIM PSNR SSIM PSNR SSIM
Ours with Fusion in Decoder 28.97 0.875 27.34 0.799 24.83 0.705
Ours w/o $w_{l}$ 30.75 0.895 27.07 0.775 26.38 0.746
Ours with Alternative Attention 30.52 0.892 27.98 0.804 26.18 0.740
Ours with $\lambda\times 5$ 29.53 0.878 27.99 0.802 26.43 0.747
Ours with $\lambda\div 5$ 29.89 0.883 27.42 0.795 26.31 0.744
Ours with Depth Anything 30.82 0.895 27.66 0.801 26.28 0.742
Ours with Depth Anything V2 30.53 0.891 27.89 0.805 26.22 0.742
Depth Feature I 30.77 0.894 26.52 0.789 26.05 0.734
Depth Feature II 30.68 0.893 27.10 0.794 26.65 0.750
NYU/KITTI fine-tune 30.75 0.895 27.69 0.802 25.68 0.728
depth estimator params+ 30.66 0.892 27.47 0.799 25.91 0.737
Ours 30.95 0.897 27.89 0.804 26.35 0.743
TABLE XIII: Results of the ablation study in the SDSD-indoor, SDSD-outdoor and SMID with SNR. Our method performs better than all ablation settings.
SDSD-indoor SDSD-outdoor SMID
Methods PSNR SSIM PSNR SSIM PSNR SSIM
Ours with Fusion in Decoder 23.50 0.877 29.14 0.836 28.50 0.806
Ours w/o $w_{l}$ 22.40 0.868 18.76 0.706 28.70 0.811
Ours with Alternative Attention 21.75 0.857 29.64 0.846 28.56 0.809
Ours with $\lambda\times 5$ 23.92 0.900 30.00 0.853 28.65 0.813
Ours with $\lambda\div 5$ 23.75 0.896 30.03 0.851 28.69 0.813
Ours with Depth Anything 27.18 0.930 28.47 0.846 28.78 0.851
Ours with Depth Anything V2 26.91 0.933 29.03 0.848 28.72 0.812
Depth Feature I 26.68 0.923 28.48 0.850 28.49 0.811
Depth Feature II 27.70 0.926 28.98 0.851 28.73 0.819
NYU/KITTI fine-tune 25.44 0.916 29.30 0.855 28.22 0.804
depth estimator params+ 25.07 0.912 28.75 0.850 28.79 0.812
Ours 26.86 0.924 29.38 0.858 29.01 0.813
TABLE XIV: Depth estimation evaluation: results of RMSE between $\hat{d}_{d}$ and $d_{n}$ in Eq 11 when employing DP3DF.
Methods SDSD-indoor SDSD-outdoor SMID
DPT 0.034 0.039 0.072
Depth Anything 0.007 0.011 0.009
Depth Anything V2 0.006 0.007 0.014
As depicted in Tables. XII and XIII, it is evident that the results of all ablation settings are inferior to our original implementation. In comparing “Ours” with “Ours with Fusion in Decoder”, we validate the impracticality of fusing restored image content features and depth-aware features since the former is close to the normal-light output while the latter is close to the low-light inputs. Furthermore, comparing “Ours” with “Ours w/o $w_{l}$.” and “Ours with Alternative Attention” underscores the significance of our proposed “Correlation-based Cross Attention for Fusion” strategy. Lastly, the consistent and even improved results across different values for $\lambda$ highlight the robustness of the hyperparameter choice in our framework. This emphasizes that researchers can choose optimal $\lambda$ values for different datasets.
Additionally, the ablation results of “Ours with Depth Anything” and “Ours with Depth Anything V2” demonstrate that our method’s performance is not dependent on the choice of the pre-trained depth estimation network. This is because, as shown in Table. XIV, our method can accurately predict depth values for low-light inputs with minimal error under the supervision provided by both DPT, Depth Anything, and Depth Anything V2.
Moreover, the comparisons among “Depth Feature I”, “Depth Feature II”, and “Ours” support that the depth-aware image features are better than depth features extracted from the depth maps in most cases. This is because depth-aware image features and low-light image features are homologous, which are more suitable for fusion. We should also note that extracting predicting depth maps directly from the low-light images will face the issue of domain shift. Also, there is a gap between the output space of the initial low-light image enhancement method and the input scope of the subsequent depth estimator, leading to depth prediction errors.
Meanwhile, comparing “NYU/KITTI fine-tune” and “Ours”, we can see that the performance with NYU and KITTI data is not superior. This may due to domain gaps between NYU/KITTI and low-light image enhancement datasets like SDSD. Furthermore, the predictions from the state-of-the-art depth estimator are generally good, not very worse than the ground truth. This is particularly true for real-world datasets like KITTI, where annotations are often sparse. Furthermore, increasing the model capability of the depth estimator may not improve the results, by comparing “depth estimator params+” and “Ours”. This might because the lack of sufficient data to train the larger model, and the ground truth employed is not changed compared with the small model.
Which is the source of performance improvement, depth information or additional feature space? In the setting of “Ours with decoder”, the fusion is conducted in the decoder part, where fusion is problematic since these two features are heterogeneous (they are close to the image and depth outputs, respectively). In this situation, although the additional feature space of the depth estimator is utilized, the performance improvement is not substantial. Thus, the improvement does not primarily stem from the addition of the feature space. On the other hand, if the fusion occurs in the encoder part, the improvement is noticeable. Consequently, depth information is a helpful source, but it requires a suitable strategy for fusion.
Figure 9: The results of the user study, which summarize that the results enhanced with our strategy are preferred by participants compared with the baselines’ results.
IV-D User Study
@@
The results of the user study are presented in Fig. 9. Clearly, the participants consistently preferred the results of our method. This indicates that our approach significantly improves the human subjective perception of existing frameworks.
IV-E Failing Cases
We provided several failure cases as shown in Fig. 10. As seen in these images, the performance of the baselines and their versions with our method are similar. These areas mainly consist of images with complex textures, where depth information is difficult to predict. As a result, the model pays less attention to utilizing the depth information in these regions. Thus, the depth priors are beneficial in areas where depth information is accurately predicted. In these regions, the depth information can provide clear boundaries between different objects, enhancing sharp boundaries and improving foreground/background separation.
We also perform several quantitative analyses to support our hypothesis. We compute the depth prediction error for each pixel location in observed failure cases by measuring the distance to the ground truth. We choose the results with DPT for experiments. As shown in Table XVI, we observe that the overall distribution of these errors is higher than the average value. This aligns with the observation that the improvement on these samples is weaker, as baseline results and those enhanced with our method are more similar, as shown in Fig. 10. Moreover, we analyze the distribution of $w_{l}$ in Eq 8 for these samples, as it represents the contribution of depth-aware features to the network backbone after the cross-attention computation. As shown in Table XVI, we find that $w_{l}$ values are lower for these failure cases, compared to other samples. These results indicate that depth-aware features are not effectively learned, and the model tends to underutilize them in these regions.
Figure 10: Failure cases. The first row shows results using SNR: the left image corresponds to SDSD-outdoor, and the right to SMID. The second row presents results using DP3DF: again, the left image is from SDSD-outdoor, and the right from SMID.
TABLE XV: Depth estimation evaluation: results of RMSE between $\hat{d}_{d}$ and $d_{n}$ in Eq 11 with DP3DF. A comparison between failure cases and other results is shown here.
Methods SDSD-outdoor SMID
DPT 0.039 0.072
Failing Case 0.127 0.089
TABLE XVI: Average value of $w_{l}$ in Eq 8 and 9 reflects the contribution of the depth-aware feature to final fused feature.
Methods SDSD-outdoor SMID
Normal Case 0.51 0.50
Failing Case 0.40 0.38
V Limitation and Future Work
@@
In this paper, we introduce a novel emphasis on employing geometrical priors, particularly in the form of depth, for low-light enhancement, an aspect overlooked by previous methods. Our extensive experiments across various datasets and networks demonstrate the effectiveness of our proposed HDGFFM. Despite achieving significant advancements, there are still areas for improvement. Firstly, further efforts are required to minimize model parameters and additional computational requirements of HDGFFM, even though it has not introduced heavy computations in this paper. Secondly, we plan to extend our strategy to incorporate geometrical priors for additional low-level vision tasks, aiming to develop a unified and generalizable theory.
Additionally, we plan to investigate the use of other types of geometric information that could potentially enhance the performance of LLE tasks. For example, we believe our method can be extended to surface normals. This is because there are state-of-the-art surface normal prediction methods that can be applied to normal-light images to obtain the ground truth.
VI Conclusion
@@
This paper presents the first attempt to leverage geometric priors (specifically depth information) for low-light enhancement, covering both image and video enhancement tasks. Rather than extracting features directly from depth maps, we propose to derive depth-aware features from low-light images, facilitating easier fusion with image features. To achieve this, we introduce a network that predicts depth information from an input low-light image. This lightweight depth estimator distills depth estimation capabilities from large-scale, open-world depth models. Furthermore, the extracted depth-aware features are efficiently fused with the encoder features of the target network using our proposed HDGFFM module, which incorporates correlation-based cross-attention to combine the strengths of both self-attention and cross-attention mechanisms. Extensive experiments on representative datasets and methods validate the effectiveness of our proposed approach.